OpenStudy (anonymous):

solve the system of linear equations using the guass jordan elimination method x - 2y =8 3x + 4y =4

OpenStudy (anonymous):

so elimination method is also the addition method so \[-2(x-2y=8)\] i will multiply by -2 so then i get -4y and the y's will cancel do you understand so far

OpenStudy (anonymous):

yes i do

OpenStudy (anonymous):

ok so you multiply by 2 not -2 sorry so will have \[2x-4y=16\] \[3x+4y=4\] so do you see now the y's cancel now solve for x

OpenStudy (anonymous):

so if you solve for x what do you have ?

OpenStudy (anonymous):

12

OpenStudy (anonymous):

\[5X=20\]

OpenStudy (anonymous):

remember elimination method is also called the addition method

OpenStudy (anonymous):

yes i remember

OpenStudy (anonymous):

:) cool hope that helped

OpenStudy (anonymous):

it did thank you very much

OpenStudy (anonymous):

gauss elimination method just like gauss jordan method